# Iterative Distillation for Reward-Guided Fine-Tuning of Diffusion Models in Biomolecular Design

### v1

@@ D.1 Details on Tasks and Reward Functions
Globularity.
Globularity refers to the extent to which a protein adopts a compact, near-spherical three-dimensional structure (Pace and Hermans, 1975). It is a hallmark of naturally occurring globular proteins, which tend to be stable, water-soluble, and functionally active. In protein generation tasks, we incorporate globularity as a reward function by minimizing the variance of the distances between each amino acid’s coordinates and the centroid, thereby encouraging the formation of highly compact and native-like structures (Hie et al., 2022).
@@ D.2 Hyperparameters
Learning rate $31e-$ 3
Reward Pred-Activity (only for fine-tuning)
@@
Learning rate $51e-$ 5
Reward (ss-match) ss-match
[…]
@@ E.1 Combining VIDD with Inference-Time Techniques
As discussed in Section 1.1, inference-time techniques are orthogonal to our method and can be combined to further enhance performances. Table 6 presents the results of applying VIDD with Best-of-N sampling (N=32), demonstrating that additional performance gains can be achieved through this combination.
Table 6: Final performance with inference-time techniques, evaluated by reward metrics and naturalness.
Method Protein SS-match Protein Globularity DNA Enhancer HepG2 Molecule Docking - Parp1
$\beta$-sheet%↑ pLDDT↑ Globularity↓ pLDDT↑ Pred-Activity↑ ATAC-Acc↑ 3-mer Corr↑ Docking Score↑ NLL↓
VIDD 0.69 $\pm$ 0.13 0.69 $\pm$ 0.11 3.96 $\pm$ 0.16 0.56 $\pm$ 0.05 7.44 $\pm$ 0.34 0.898 $\pm$ 0.302 0.509 9.4 $\pm$ 1.7 741 $\pm$ 21
VIDD + BoN 0.79 $\pm$ 0.01 0.71 $\pm$ 0.07 3.77 $\pm$ 0.06 0.61 $\pm$ 0.04 7.63 $\pm$ 0.33 0.813 $\pm$ 0.390 0.507 12.1 $\pm$ 1.0 726 $\pm$ 28
E.2 Futher Results for Molecule Generation
First, we report the diversity comparisons across methods in Table 7 on page 7.
@@ E.2 Futher Results for Molecule Generation
To evaluate the validity of our method in molecule generation, we further report several key metrics that capture different aspects of molecule quality and diversity in Table 8 on page 8.
@@
We compare the metrics using 512 molecules generated from the pre-trained GDSS model and from different methods, as shown in Table 8 on page 8. Overall, our method achieves comparable performances with the pre-trained model on all metrics, maintaining high validity, novelty, and uniqueness while outperforming on several metrics such as FCD, SNN, and NSPDK MMD. Pre-trained performs consistently well across all metrics, particularly in SNN and atomic stability. However, it does not optimize specific molecular properties as effectively as the other methods. DDPP performs poorly in scaffold similarity and NSPDK MMD, indicating that it generates unrealistic molecules. These results indicate that our approach can generate a diverse set of novel molecules that are chemically plausible and relevant.
E.3 Further Results for Protein Generation
Table 9 presents the protein generation results evaluated with diversity metrics on ss-match and globularity optimization tasks. Diversity is computed based on token-level differences among a batch of generated sequences (batch size = 32). Note that diversity serves as a secondary metric compared to reward quality since different random seeds can yield diverse outputs as long as the model does not collapse in practice. As shown in the table, all methods maintain a reasonable level of diversity, indicating no signs of mode collapse.
@@ E.3 Further Results for Protein Generation
Table 12 presents results obtained when training is continued until mode collapse, where the model produces highly similar sequences with negligible diversity. Although this setting achieves higher reward values, it reflects poor generative quality. We report these results to highlight the trade-off between reward maximization and maintaining diversity during training.
Table 9: Performance of different methods on protein generation w.r.t. reward, pLDDT, and diversity.
Method Protein SS-match Protein Globularity
$\beta$-sheet%↑ pLDDT↑ Diversity↑ Globularity↓ pLDDT↑ Diversity↑
Pre-trained 0.05 $\pm$ 0.05 0.37 $\pm$ 0.09 0.908 7.42 $\pm$ 2.59 0.37 $\pm$ 0.09 0.908
Best-of-N (N=32) 0.26 $\pm$ 0.13 0.38 $\pm$ 0.11 0.904 5.05 $\pm$ 0.21 0.39 $\pm$ 0.14 0.905
Standard Fine-tuning 0.11 $\pm$ 0.12 0.32 $\pm$ 0.06 0.436 6.30 $\pm$ 1.11 0.31 $\pm$ 0.04 0.907
DDPP 0.60 $\pm$ 0.07 0.40 $\pm$ 0.08 0.898 5.98 $\pm$ 0.89 0.32 $\pm$ 0.10 0.902
DDPO 0.19 $\pm$ 0.17 0.35 $\pm$ 0.10 0.908 4.75 $\pm$ 0.20 0.37 $\pm$ 0.06 0.904
VIDD 0.69 $\pm$ 0.13 0.69 $\pm$ 0.11 0.852 3.96 $\pm$ 0.16 0.56 $\pm$ 0.05 0.757
Table 10: Performance of different methods on protein binding design tasks on target protein PD-L1. To prevent model collapse, we report results with diversity $\geq$ 0.5.
@@
Table 11: Performance of different methods on protein binding design tasks on target protein IFNAR2. To prevent model collapse, we report results with diversity $\geq$ 0.5.
@@
Table 12: Performance comparison on protein binding design tasks without diversity constraints. The results show the trade-off between reward maximization and diversity.
Method Reward↑ ipTM↑ pLDDT↑ Radius↓ Diversity↑
PD-L1
VIDD (collapsed) 0.9559 $\pm$ 0.0002 0.8571 $\pm$ 0.0002 0.9411 $\pm$ 0.0001 -0.2358 $\pm$ 0.0001 0.0006
IFNAR2
VIDD (collapsed) 0.8277 $\pm$ 0.0004 0.7706 $\pm$ 0.0002 0.7079 $\pm$ 0.0017 0.6829 $\pm$ 0.0017 0.0006
@@ Appendix F Visualization of Generated Samples
Figure 7: Visualization of generated molecules using VIDD optimizing the reward of docking score for parp1 (normalized as $max(-DS,0)$).

### v2

@@ D.2 Hyperparameters
Learning rate 1e41e-4
Reward Pred-Activity
@@
Learning rate 1e51e-5
Reward (ss-match) $s$-match
@@ E.1 Combining VIDD with Inference-Time Techniques
As discussed in Section 1.1, inference-time techniques are orthogonal to our method and can be combined to further enhance performances. Table 6 and Table 7 presents the results of applying VIDD with Best-of-N sampling, demonstrating that additional performance gains can be achieved through this combination.
Table 6: Final performance combined with inference-time techniques.
Method Protein SS-match DNA Enhancer HepG2 Molecule Docking - Parp1
$\beta$-sheet%↑ pLDDT↑ Pred-Activity↑ ATAC-Acc↑ 3-mer Corr↑ Docking Score↑ NLL↓
VIDD 0.83 $\pm$ 0.01 0.82 $\pm$ 0.01 8.28 $\pm$ 0.18 0.820 $\pm$ 0.384 0.162 9.4 $\pm$ 1.7 741 $\pm$ 21
VIDD + BoN (N=32) 0.84 $\pm$ 0.00 0.82 $\pm$ 0.01 8.40 $\pm$ 0.07 0.750 $\pm$ 0.433 0.152 12.1 $\pm$ 1.0 726 $\pm$ 28
Table 7: Final performance combined with inference-time techniques for protein binder design.
Method PD-L1 IFNAR2
ipTM↑ pLDDT↑ Radius↓ Diversity↑ ipTM↑ pLDDT↑ Radius↓ Diversity↑
VIDD 0.82 $\pm$ 0.02 0.87 $\pm$ 0.04 -0.12 $\pm$ 0.11 0.55 0.51 $\pm$ 0.11 0.47 $\pm$ 0.05 2.20 $\pm$ 2.00 0.52
VIDD + BoN (N=128) 0.84 $\pm$ 0.00 0.91 $\pm$ 0.01 -0.16 $\pm$ 0.06 0.39 0.66 $\pm$ 0.02 0.55 $\pm$ 0.04 1.04 $\pm$ 0.45 0.43
Table 8: The influence of lazy update interval $K$ on the performances of DNA sequence design.
[…]
1 7.06 $\pm$ 0.35 0.000 $\pm$ 0.000 0.211
5 8.28 $\pm$ 0.18 0.820 $\pm$ 0.384 0.162
10 7.76 $\pm$ 0.32 0.047 $\pm$ 0.211 0.457
20 7.71 $\pm$ 0.37 0.086 $\pm$ 0.280 0.398
E.2 Further Results for DNA Generation
We study the effect of the lazy update interval $K$ on the DNA sequence design. As shown in Table 8, performance is different under different $K$ and does not improve monotonically with more frequent updates. These results suggest that less updates of roll-out policy can stabilize training and improve optimization.
E.3 Further Results for Molecule Generation
First, we report the diversity comparisons across methods in Table 9 on page 9.
@@ E.3 Further Results for Molecule Generation
To evaluate the validity of our method in molecule generation, we further report several key metrics that capture different aspects of molecule quality and diversity in Table 10 on page 10.
@@
We compare the metrics using 512 molecules generated from the pre-trained GDSS model and from different methods, as shown in Table 10 on page 10. Overall, our method achieves comparable performances with the pre-trained model on all metrics, maintaining high validity, novelty, and uniqueness while outperforming on several metrics such as FCD, SNN, and NSPDK MMD. Pre-trained performs consistently well across all metrics, particularly in SNN and atomic stability. However, it does not optimize specific molecular properties as effectively as the other methods. DDPP performs poorly in scaffold similarity and NSPDK MMD, indicating that it generates unrealistic molecules. These results indicate that our approach can generate a diverse set of novel molecules that are chemically plausible and relevant.
E.4 Further Results for Protein Generation
More metrics on protein binder design.
@@ E.4 Further Results for Protein Generation
Table 11: Performance of different methods on protein binding design tasks on target protein PD-L1.
@@
Table 12: Performance of different methods on protein binding design tasks on target protein IFNAR2.
@@
Figure 5: Performance vs. diversity in PD-L1 binder design under different roll-in mixtures. Mixing in the pre-trained policy during roll-in (smaller $\beta_{s}$) increases diversity compared to relying solely on the roll-out policy ($\beta_{s}{=}1$).
Mixed roll-in policy
In Section 4.1, we describe a mixed roll-in strategy that samples trajectories from the pre-trained policy $p^{\text{pre}}_{t}$ and the roll-out policy $p^{\mathrm{out}}_{t}$ with probabilities $1-\beta_{s}$ and $\beta_{s}$, respectively. Figure 5 ablates $\beta_{s}$ and find that injecting a non-zero fraction of $p^{\text{pre}}_{t}$ (i.e., not always rolling in from $p^{\mathrm{out}}_{t}$) improves diversity and often yields better overall performance. The specific optimal mixture is task-dependent; accordingly, we treat $\beta_{s}$ as a tunable hyperparameter selected by validation to balance exploration ($p^{\text{pre}}_{t}$) and exploitation ($p^{\mathrm{out}}_{t}$).
(a) $S$-match
(b) PD-L1
(c) IFNAR2
Figure 6: Training curves of different methods on $S$-match, PD-L1, and IFNAR2 binder design tasks. The $y$-axis shows the optimized reward, and the $x$-axis shows training steps.
(a) $S$-match
(b) PD-L1
(c) IFNAR2
Figure 7: Scatter plots of reward versus diversity for different methods on $S$-match, PD-L1, and IFNAR2 binder design tasks. Each point corresponds to a training checkpoint.
Trade-off between Reward and Diversity
Readers may be concerned that higher rewards could come at the cost of reduced diversity from results in Table 1 and Table 2 (as well as Table 11 and Table 12). To clarify this, we include the training curves (Figure 6) and the reward–diversity scatter plots (Figure 7). The curves show how reward evolves over training, while the scatter plots visualize the reachable regions in the reward–diversity plane for each method. These figures illustrate that only our method is able to expand the reachable set and achieve substantially higher rewards across training.
@@ Appendix F Visualization of Generated Samples
Figure 8: Visualization of generated molecules using VIDD optimizing the reward of docking score for parp1 (normalized as max(DS,0)max(-DS,0)).